OpenStudy (anonymous):

A speeding ticket costs $40 plus $5 for each mile per hour over the speed limit. Lynne mentions that she was given a ticket for $75. What is a valid conclusion using deductive reasoning?

OpenStudy (anonymous):

40 + 5x = 75 5x = 75 - 40 = 35 x = 35/5 = 7 Lynne was 7 miles over the speed limit.

OpenStudy (anonymous):

40+ 5x=75